OpenStudy (anonymous):

Sorry about that. i was trying to upgrade to IE9 but it would not allow me to sign in. So I went thru all of last night and did not get the 80% I need to continue on. So if you don't mind please just give me the answers so I can move on. Thank you in advance. Complete the ordered pair below for the equation y=-3x+2. (0,__)

OpenStudy (anonymous):

Plug in 0 for x and solve for y.

OpenStudy (anonymous):

You won't be able to understand it, if you don't figure it out yourself.

OpenStudy (anonymous):

It really isn't that hard of a problem if you just try.

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

well like I said I just don't have the time and need to get thru this

OpenStudy (anonymous):

y=-3(0)+2 y=0+2 y=2 (0,2) Hope this helps! :)

OpenStudy (anonymous):

thanks spent hours and hours yesterday and just missed passing it

OpenStudy (anonymous):

no problemmm